# Sparse Autoencoders for Scientifically Rigorous Interpretation of Vision Models

### v1

@@ 4 SAE-Enabled Analysis of Vision Models
Sparse autoencoders (SAEs) provide a powerful new lens for understanding and comparing vision models. By decomposing dense activation vectors into interpretable features, SAEs enable systematic analysis of what different architectures learn and how their training objectives shape their internal representations. We demonstrate that even simple manual inspection of top-activating images for SAE-discovered features can reveal fundamental differences between models that would be difficult to detect through other methods.
@@
While prior work has used techniques like TCAV (Kim et al., 2018) to probe for semantic concepts in vision models, these methods typically test for pre-defined concepts and rely on supervised concept datasets. In contrast, our SAE-based approach discovers interpretable features directly from model activations without requiring concept supervision. We train SAEs on intermediate layer activations following the procedure detailed in Section 3, then analyze the highest-activating image patches for each learned feature as in Figure 3. This straightforward process reveals both specific features (e.g., a feature that fires on dental imagery) and model-level patterns (e.g., one model consistently learning more abstract features than another).
@@
We refer to individual SAE features using a standardized format MODEL-WIDTH/INDEX, where MODEL identifies the vision transformer the SAE was trained on (e.g., CLIP or DINOv2), WIDTH indicates the number of features in the SAE (e.g., 24K for $24,576$), and INDEX uniquely identifies the specific feature. For example, CLIP-24K/20652 refers to feature $20,652$ from a $24,576$-dimensional SAE trained on CLIP activations.
4.1 Language Supervision Enables Cultural Understanding
@@
We analyze SAE features and find that CLIP learns remarkably robust representations of cultural and geographic concepts–a capability that appears absent in DINOv2’s purely visual representations. This finding demonstrates how language supervision enables learning of abstract cultural features that persist across diverse visual manifestations.
@@
Consider CLIP’s representation of country-specific visual features. We find individual SAE features that consistently activate on imagery associated with specific countries, while remaining inactive on visually similar but culturally distinct images (Figure 4). For instance, CLIP reliably detects German visual elements across architectural landmarks (like the Brandenburg Gate), sports imagery (German national team uniforms), and cultural symbols (Oktoberfest celebrations). Crucially, this feature remains inactive on other European architectural landmarks or sporting events, suggesting it captures genuine cultural associations rather than just visual similarities.
@@
Similarly, we find features that activate on distinctly Brazilian imagery, spanning Rio de Janeiro’s urban landscape, the national flag, coastal scenes, and cultural celebrations. These features show selective activation, responding strongly to Brazilian content while remaining inactive on visually similar scenes from other South American locations. This selective response pattern suggests CLIP has learned to recognize and group culturally related visual elements, even when they share few low-level visual features. See Figure B1 for additional examples of this phenomena.
@@
In contrast, when we analyze DINOv2’s features, we find no comparable country-specific representations. This fundamental difference reveals how language supervision guides CLIP to learn culturally meaningful visual abstractions that pure visual training does not discover.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
4.2 Language Supervision Induces Semantic Abstraction
Beyond cultural concepts, CLIP learns abstract semantic features that persist across visual styles–a capability absent in DINOv2–revealing how language supervision enables human-like semantic concept formation.
For example, we discover an SAE feature (CLIP-24K/20652) that activates on accident scenes across photographs of car accidents, crashed planes and cartoon depictions of crashes (Figure 5). This feature fires strongly whether processing a news photograph of a car accident or a stylized illustration of a collision–suggesting CLIP has learned a truly abstract representation of “accident” that transcends visual style.
In contrast, DINOv2’s features fragment these examples across multiple low-level visual patterns. While it learns features that detect specific visual aspects of accidents (like crumpled metal in photographs), no single feature captures the semantic concept across different visual styles. This suggests that without language supervision, DINOv2 lacks the learning signal needed to unite these diverse visual presentations into a single abstract concept. We hypothesize that CLIP’s language supervision provides explicit signals to group visually distinct but semantically related images, while DINOv2’s purely visual training offers no such bridge across visual styles. The ability to form such abstract semantic concepts, independent of visual style, represents a fundamental difference in how these models process and represent visual information.
@@ 4.3 Implications for Vision Model Selection
Recent work has demonstrated empirical benefits from combining different vision encoders in vision-language models (Liu et al., 2024b; a; Lu et al., 2024). Tong et al. (2024b) constructed a challenging dataset for VLMs using only CLIP as a visual encoder. Cambrian-1 (Tong et al., 2024a) found improved performance when incorporating both CLIP and DINOv2 encoders, and Jiang et al. (2023) systematically examined the distinct contributions of different visual encoders in VLMs. However, the mechanistic basis for these improvements has remained unclear.
@@
Figure 6: Demonstrating the scientific method for understanding vision model behavior using sparse autoencoders (SAEs). Left: We observe that CLIP predicts “Blue Jay.” Upper Middle: We select the bird’s wing in the input image; the SAE proposes a hypothesis that the most salient feature is “blue feathers” via exemplar images. Lower Middle: We validate this hypothesis through controlled intervention by suppressing the identified “blue feathers” feature in the model’s activation space. Right: we observe a change in behavior: the predicted class shifts away from “Blue Jay” towards “Clark Nutcracker”, a similar bird besides the lack of blue plumage. This three-step process of observation, hypothesis formation, and experimental validation enables systematic investigation of how vision models process visual information.
5 Validating Hypotheses of Vision Model Behavior
Understanding how vision models process information requires proving that our explanations accurately capture the model’s behavior. While prior work has demonstrated interpretable features or model control in isolation, validating explanations requires showing both that discovered features are meaningful and that we can precisely manipulate them. We demonstrate that SAE-derived features enable reliable control across two complementary visual understanding tasks–classification and segmentation–providing strong evidence that our interpretations faithfully capture model behavior.
Vision models process information across multiple levels of abstraction, from local visual features to global semantic reasoning. Standard evaluation approaches focus on a single level of processing, limiting their ability to validate feature explanations. We overcome this by strategically selecting tasks that test different aspects of visual understanding: classification validates precise control over visual attributes and segmentation demonstrates spatial coherence of identified features,
For each task, we first train or utilize an existing task-specific prediction head. We then perform controlled feature interventions using our SAE and compare model behavior before and after intervention. Success across these diverse challenges, which we demonstrate through extensive qualitative results, provides strong evidence that our method identifies meaningful and faithful features.
@@ 5.1 Image Classification
Precisely manipulating individual visual features is essential for validating our understanding of how vision models make decisions. While prior work identifies features through post-hoc analysis, proving causal relationships requires demonstrating that controlled feature manipulation yields predictable output changes. Using fine-grained bird classification as a testbed, we show that SAE-derived features enable precise control: we can manipulate specific visual attributes like beak shape or plumage color while preserving other traits, producing predictable changes in classification outputs that validate our feature-level explanations of model behavior.
We train a sparse autoencoder on activations from layer $11$ of a CLIP-pretrained ViT-B/16 (Radford et al., 2021), using the complete ImageNet-1K dataset (Russakovsky et al., 2015) to ensure broad feature coverage. The SAE uses a 32×32$\times$32 × expansion factor ($24,576$ features) to capture a rich vocabulary of visual concepts. Through an interactive exploration interface, we identify interpretable features by examining patches that maximally activate specific SAE features. For example, selecting the blue feathers of a blue jay (Cyanocitta cristata) reveals SAE features that consistently activate on similar colorations across the dataset.
To validate the proposed explanation of these features, we can manipulate them by adjusting their values in the SAE’s latent space. When we reduce the “blue feathers” feature in a blue jay image, the model’s prediction shifts from blue jay to Clark’s nutcracker (Nucifraga columbiana)–a semantically meaningful change that aligns with ornithological knowledge (Figure 6). See Appendix C for SAE details, linear classifier details, and additional qualitative examples.
Figure 7: Far Left: We train a linear head to predict semantic segmentation classes for each patch. Middle Left: We choose all sand-filled patches in the input image to inspect. Middle Right: Our SAE proposes exemplar images for the maximally activating sparse dimensions, as in Section 5.1, suggesting that DINOv2 is learning a sand feature. Far Right: We suppress the sand feature in not just the selected patches, but all patches. We modify all activation vectors and pass them to DINOv2’s final transformer layer followed by our trained linear segmentation head. We see that the head predicts “earth, ground” and “water” for the former sand patches. Both classes are good second choices if “sand” is unavailable. Notably, other patches are not meaningfully affected, demonstrating the pseudo-orthogonality of the SAE’s learned feature vectors.
@@ 5.2 Semantic Segmentation
Moving beyond single-feature control requires proving that discovered features can be manipulated independently—a key challenge that traditional interpretability methods struggle to address. When features are entangled, attempts to modify one aspect of an image often produce unintended changes in others, making precise intervention impossible. Through semantic segmentation, we demonstrate that SAE features form a pseudo-orthogonal basis for the model’s representational space—while not mathematically perpendicular, these features are functionally independent in practice. When we suppress semantic concepts like “sand” or “grass”, we observe consistent changes in targeted regions while leaving other predictions intact, demonstrating this functional independence.
We train a linear probe on vision model patch-level features on the ADE20K semantic segmentation dataset (Zhou et al., 2017) following the setup in Oquab et al. (2023). Specifically, given an image-level representation tensor $\mathbb{R}^{p\times p\times d}$, we train a linear probe to predict a low-resolution $p\times p\times C$ logit map, which is up-sampled to full resolution. We achieve $35.1$ mIoU on the ADE20K validation set. While this method falls short of state-of-the-art methods, it is a sufficient testbed for studying feature control. Specific training details are are available in Appendix D.
We train a sparse autoencoder on ImageNet-1K activations from layer 11 of a DINOv2-pretrained ViT-B/16 following the procedure described in Section 3.
To validate our method’s ability to manipulate semantic features, we developed an interactive interface that allows precise control over individual SAE features. Users can select specific image patches and modify their feature activations in both positive and negative directions, enabling targeted manipulation of semantic concepts. This approach provides several advantages over automated evaluation metrics: it allows exploration of feature interactions, demonstrates the spatial coherence of manipulations, and reveals how features compose to form higher-level concepts. The interface makes it possible to test hypotheses about learned features through direct experimentation–for instance, we can verify that “sand” features truly capture sand-like textures by suppressing them and observing consistent changes across diverse images, as shown in Figure 7.
@@ 6 Conclusion
We demonstrate that sparse autoencoders enable both hypothesis formation and controlled testing for vision models, enabling both interpretation of learned features and precise control over model behavior.
Our work reveals fundamental insights about representation learning in vision models. We show that language supervision guides CLIP toward human-interpretable abstractions that generalize across visual styles, while purely visual training in DINOv2 produces more granular, style-specific features. This finding suggests that achieving human-like visual abstraction may require bridging between different modalities or forms of supervision.
Through controlled experiments across classification, segmentation, and image-text tasks, we validate that SAE-discovered features capture genuine causal relationships in model behavior rather than mere correlations. The ability to reliably manipulate these features while maintaining semantic coherence demonstrates that vision transformers learn decomposable, interpretable representations even without explicit supervision.
However, significant challenges remain. Current methods for identifying manipulable features still require manual exploration, and the relationship between feature interventions and model behavior becomes increasingly complex for higher-level tasks. Future work should focus on automating feature discovery, understanding feature interactions across different model architectures, and extending reliable control to more sophisticated visual reasoning tasks.
More broadly, our results suggest that combining interpretation with control may be essential for developing truly trustworthy AI systems. Just as scientific understanding emerges from our ability to both explain and manipulate natural phenomena, meaningful understanding of neural networks requires tools that unite explanation with experimentation. We believe SAEs provide a promising foundation for building such tools.
Acknowledgment
We thank colleagues from the Imageomics Institute and the OSU NLP group for their constructive feedback. This research is supported by National Science Foundation OAC-2118240. We are thankful for the generous support of the computational resources by the Ohio Supercomputer Center.
@@ Acknowledgment
    Appendix A: SAE Training Details (for Section 3.2)
@@
    Appendix B: Additional Examples (for Section 4)
@@
    Appendix C: Classification Details (for Section 5.1)
@@
    Appendix D: Classification Details (for Section 5.2)
Appendix A SAE Training Details

### v2

@@ 4 SAE-Enabled Analysis of Vision Models
Sparse autoencoders (SAEs) provide a useful new lens for understanding and comparing vision models. By decomposing dense activation vectors into interpretable features, SAEs enable systematic analysis of what different architectures learn and how their training objectives shape their internal representations. We demonstrate that even simple manual inspection of top-activating images for SAE-discovered features can reveal differences between models that would be difficult to detect through other methods.
@@
While prior work has used techniques like TCAV [kim2018tcav] to probe for semantic concepts in vision models, these methods typically test for pre-defined concepts and rely on supervised concept datasets. In contrast, our SAE-based approach discovers interpretable features directly from model activations without requiring concept supervision. We train SAEs on intermediate layer activations following the procedure detailed in Section˜3, then analyze the highest-activating image patches for each learned feature as in Fig.˜2. This straightforward process reveals both specific features (e.g., a feature that fires on dental imagery) and model-level patterns (e.g., one model consistently learning more abstract features than another). Technical details of our process are provided in Section˜9.
@@
We refer to individual SAE features using MODEL-WIDTH/INDEX, where MODEL identifies the vision transformer the SAE was trained on (e.g., CLIP or DINOv2), WIDTH indicates the number of features in the SAE (e.g., 24K for $24,576$), and INDEX uniquely identifies the specific feature. For example, CLIP-24K/20652 refers to feature $20,652$ from a $24,576$-dimensional SAE trained on CLIP activations.
Table 1: SAE metrics on ImageNet-1K. Low reconstruction error (mean-squared error; MSE) and sparse activations demonstrate successful decomposition of ViT representations. “Dead” neurons are active on less than $10^{-7}\%$ of inputs; “Dense” neurons are active on more than $1\%$ of all inputs.
Model MSE L0 Dead Dense
CLIP 0.0761 412.7 0 11,858
DINOv2 0.0697 728.7 1 19,735
Figure 4: MSE and L0 on the training dataset for all learning rates and sparsity coefficents $\lambda$.
4.1 Language Enables Cultural Understanding
@@
We analyze SAE features and find that CLIP learns remarkably robust representations of cultural and geographic concepts–a capability that appears absent in DINOv2’s purely visual representations. This demonstrates how language supervision enables learning abstract cultural features that persist across diverse visual manifestations.
@@
We find individual SAE features that consistently activate on imagery associated with specific countries, while remaining inactive on visually similar but culturally distinct images (Fig.˜3). For instance, CLIP reliably detects German visual elements across architectural landmarks (like the Brandenburg Gate), sports imagery (German national team uniforms), and cultural symbols (Oktoberfest celebrations). Crucially, this feature remains inactive on other European architectural landmarks or sporting events, suggesting it captures genuine cultural associations rather than just visual similarities.
@@
Similarly, we find features that activate on distinctly Brazilian imagery, spanning Rio de Janeiro’s urban landscape, the national flag, coastal scenes, and cultural celebrations. These features show selective activation, responding strongly to Brazilian content while remaining inactive on visually similar scenes from other South American locations. This selective response pattern suggests CLIP has learned to recognize and group culturally related visual elements, even when they share few low-level visual features. See Fig.˜8 for additional examples of this phenomena.
@@
In contrast, when we analyze DINOv2’s features, we find no comparable country-specific representations. This difference illustrates how language supervision could help CLIP to learn culturally meaningful visual abstractions that pure visual training does not discover.
@@
4.2 Language Induces Semantic Abstraction
Beyond cultural concepts, CLIP learns abstract semantic features that persist across visual styles, a capability DINOv2 lacks, revealing how language supervision enables human-like semantic learning.
A striking example emerges in representations of mechanical accidents and crashes. We discover a feature (CLIP-24K/20652) that activates on accident scenes across photographs of car accidents, crashed planes and cartoon depictions of crashes (Fig.˜5). This feature fires on both a news photograph of a car accident or a stylized illustration of a collision–suggesting CLIP learns an abstract representation of “accident” that transcends visual style.
In contrast, DINOv2’s features fragment these examples across multiple low-level visual patterns. While it learns features that detect specific visual aspects of accidents (like crumpled metal in photographs), no single feature captures the semantic concept across different visual styles. This suggests that without language supervision, DINOv2 lacks the learning signal needed to unite these diverse visual presentations into a single abstract concept. We hypothesize that CLIP’s language supervision provides explicit signals to group visually distinct but semantically related images, while DINOv2’s purely visual training offers no such bridge across visual styles. The ability to form such abstract semantic concepts, independent of visual style, is a meaningful difference in how these models process and represent visual information.
@@ 4.3 Implications for Vision Model Selection
[tong2024eyes] constructed a challenging dataset for vision-language models (VLMs; [liu2024llava, liu2024llava1_5, lu2024deepseek_vl]) using only CLIP as a visual encoder. Prior work [jiang2023clip, shi2024eagle, tong2024cambrian, li2025eagle] demonstrates empirical benefits from combining different vision encoders. However, the mechanistic basis for these improvements has remained unclear.
@@
Figure 6: Demonstrating the scientific method for understanding vision model behavior using sparse autoencoders (SAEs). Left: We observe that CLIP predicts “Blue Jay.” Upper Middle: We select the bird’s wing in the input image; the SAE proposes a hypothesis that the most salient feature is “blue feathers” via exemplar images. Lower Middle: We validate this hypothesis through controlled intervention by suppressing the identified “blue feathers” feature in the model’s activation space. Right: we observe a change in behavior: the predicted class shifts away from “Blue Jay” towards “Clark Nutcracker”, a similar bird besides the lack of blue plumage. This three-step process of observation, hypothesis formation, and validation enables systematic investigation of how vision models process visual information.
5 Testing Hypotheses of Vision Model Behavior
Understanding how vision models process information requires testing that our explanations accurately capture the model’s behavior. While prior work has demonstrated interpretable features or model control in isolation, validating explanations requires showing both that discovered features are meaningful and that we can precisely manipulate them. We demonstrate that SAE-derived features enable reliable control across two complementary visual understanding tasks–classification and segmentation–providing strong evidence that our interpretations faithfully capture model behavior.
For each task, we first train or utilize an existing task-specific prediction head. We then perform controlled feature interventions using our SAE and compare model behavior before and after intervention. Success across these diverse challenges, which we demonstrate through diverse qualitative results, provides strong evidence that our method identifies meaningful and faithful features.
@@ 5.1 Image Classification
Precisely manipulating individual visual features is essential for validating our understanding of vision model decisions. While prior work identifies features through post-hoc analysis, testing causal relationships requires demonstrating that controlled feature manipulation yields predictable output changes. Using fine-grained bird classification as a testbed, we show that SAE-derived features enable precise control: we can manipulate specific visual attributes like beak shape or plumage color while preserving other traits, producing predictable changes in classification outputs that validate our feature-level explanations of model behavior.
We train a sparse autoencoder on activations from a CLIP-pretrained ViT-B/16 [radford2021learning], using the complete ImageNet-1K dataset [russakovsky2015imagenet] to ensure broad feature coverage. We record activations from layer $11$ (of $12$ total layers); we argue that the second-to-last layer contains features that are highly semantic but not overfit to CLIP’s contrastive image-text matching objective. The SAE uses a $32\times$ expansion factor ($24,576$ features) to capture a rich vocabulary of visual concepts. Through an interactive interface, we identify interpretable features by examining patches that maximally activate specific SAE features. For example, selecting the blue feathers of a blue jay (Cyanocitta cristata) reveals SAE features that consistently activate on similar colorations across the dataset.
To validate the proposed feature explanation, we can manipulate them by adjusting their values in the SAE’s latent space. When we reduce the “blue feathers” feature in a blue jay image, the model’s prediction shifts from blue jay to Clark’s nutcracker (Nucifraga columbiana)–a semantically meaningful change that aligns with ornithological knowledge (Fig.˜6). See Section˜10 for SAE details, linear classifier details, and additional qualitative examples. While our current interface limits manipulation to single features at a time, the approach generalizes beyond bird classification and provides a useful new tool for understanding vision model decisions.
Figure 7: Far Left: We train a linear head to predict semantic segmentation classes for each patch. Middle Left: We choose all “sand” patches in the input image to inspect. Middle Right: Our SAE proposes exemplar images for the maximally activating sparse dimension, as in Section˜5.1, suggesting that DINOv2 learns a sand feature. Far Right: We suppress the sand feature in not just the selected patches, but all patches. We modify all activation vectors before DINOv2’s final transformer layer followed by our trained linear segmentation head. We see that the head predicts “earth, ground” and “water” for the former sand patches. Both classes are good second choices if “sand” is unavailable. Notably, other patches are not meaningfully affected, demonstrating the pseudo-orthogonality of the SAE’s learned feature vectors.
@@ 5.2 Semantic Segmentation
When discovered features are entangled, attempts to modify one aspect of an image often produce unintended changes in others, making precise intervention impossible. Through semantic segmentation, we observe that SAE features behave as a pseudo-orthogonal basis for the model’s representational space; while not mathematically perpendicular, these features are, in practice, functionally independent. When we suppress semantic concepts like “sand” or “grass” in all patches, we observe consistent changes in targeted regions while leaving other predictions intact, demonstrating this functional independence.
We train a linear probe on vision model patch-level features on the ADE20K semantic segmentation dataset [zhou2017ade20k] following the setup in [oquab2023dinov2]. Specifically, given an image-level representation tensor of $p\times p$ patch vectors of dimension $d$ ($\mathbb{R}^{p\times p\times d}$), we train a linear probe to predict a low-resolution $p\times p\times C$ logit map, which is up-sampled to full resolution. We achieve $35.1$ mIoU on the ADE20K validation set. While this method falls short of state-of-the-art methods, it is a sufficient testbed for studying feature control. We train a sparse autoencoder on ImageNet-1K activations from layer 11 of a DINOv2-pretrained ViT-B/16 following the procedure described in Section˜3. We switch from CLIP to DINOv2 to demonstrate that our experimental validation approach generalizes across model families with distinct training objectives and because DINOv2 provides richer spatial representations for dense prediction tasks like semantic segmentation. Details are are available in Section˜11.
To validate our method’s ability to manipulate semantic features, we developed an interactive interface that allows precise control over individual SAE features. Users can select specific image patches and modify their feature activations in both positive and negative directions, enabling targeted manipulation of semantic concepts. This approach provides several advantages over automated evaluation metrics: it allows exploration of feature interactions, demonstrates the spatial coherence of manipulations, and reveals how features compose to form higher-level concepts. The interface makes it possible to test hypotheses about learned features through direct experimentation; for example, we can verify that “sand” features truly capture sand-like textures by suppressing them and observing consistent changes across diverse images, as shown in Fig.˜7.
@@ 6 Conclusion
We demonstrate that sparse autoencoders enable both hypothesis formation and controlled testing for vision models, enabling empirically validated model interpretation.
Our work reveals fundamental insights about representation learning in vision models. Our qualitative study suggests that language supervision helps CLIP develop abstractions that generalize across visual styles, a pattern absent in our observations of DINOv2. This finding suggests that achieving human-like visual abstraction may require bridging between different modalities or forms of supervision.
We find that SAE-discovered features reliably capture genuine causal relationships in model behavior rather than mere correlations in classification and semantic segmentation. The ability to reliably manipulate these features while maintaining semantic coherence demonstrates that vision transformers learn decomposable, interpretable representations even without explicit supervision.
However, significant challenges remain. Current methods for identifying manipulable features still require manual exploration, and the relationship between feature interventions and model behavior becomes increasingly complex for higher-level tasks (see Section˜7 for additional discussion of limitations). Future work should focus on automating feature discovery, understanding feature interactions across different model architectures, and extending reliable control to more sophisticated visual reasoning tasks.
Just as scientific understanding emerges from our ability to both explain and manipulate natural phenomena, meaningful understanding of neural networks requires tools that unite explanation with experimentation. We believe SAEs provide a promising foundation for building such tools.
@@
    Section˜7: Limitations
@@
    Section˜8: SAE Training Details (for Section˜3.2)
@@
    Section˜9: SAE-Enabled Analysis Details (for Section˜4)
@@
    Section˜10: Classification Details (for Section˜5.1)
  5. 5.
    Section˜11: Semantic Segmentation Details (for Section˜5.2)
  6. 6.
    Section˜12: Compute Requirements
7 Limitations
Manual, potentially selective feature mining.
Discovering interesting SAE units still depends on human browsing of exemplar grids. This process is slow, may overlook important concepts, and risks cherry-picking—an issue flagged by reviewers. Automated search and concept-label assignment (e.g. via language supervision) remain important future directions.
Qualitative evidence only.
Our work relies significantly on qualitative evaluation through visual inspection of feature activations and manipulation effects. While this approach provides intuitive validation of feature interpretations, it lacks standardized quantitative metrics for measuring both interpretability and control effectiveness. Developing robust metrics that quantify semantic alignment, manipulation specificity, and cross-model comparability would advance this research direction from exploratory investigation toward systematic evaluation. Such metrics would facilitate comparing different interpretability methods and tracking progress in the field.
Causality caveats and distribution shift.
Latent edits approximate knockout experiments, yet large interventions can push activations off-manifold. Feature correlations and non-linear interactions complicate clean causal attribution, limiting the strength of causal conclusions.
Coverage and stability of the concept dictionary.
SAEs are not guaranteed to surface every concept a model encodes [fel2025archetypal]. Low-activation units can be polysemantic and may vary across random initialisations. Systematic studies of exhaustiveness and retraining stability are deferred to future work.
Task scope and baseline overlap.
We validate only on fine-grained classification and semantic segmentation. Other tasks (e.g. bias removal, captioning) and prior frameworks such as DN-CBM already enable editable concepts for classification; broader task coverage and explicit baselines would clarify added value.
Pitfalls of quantitative faithfulness metrics.
Prior work shows that standard saliency “faithfulness scores” can be misleading under distribution shift [hooker2019benchmark, adebayo2018sanity]. Following the qualitative-research guidance of [olah2024qualitative], we prioritise human-falsifiable dashboards until community-accepted metrics emerge.
8 SAE Training Details